### v1

# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

Abstract.
Coarse-grained reconfigurable arrays (CGRAs), or dataflow architectures in general, have gained attention in recent years due to their promising power efficiency compared to traditional von Neumann architectures. To program these architectures using ordinary languages such as C, a dataflow compiler must transform the original sequential, imperative program into an equivalent dataflow graph, composed of dataflow operators running in parallel. This transformation is challenging since the asynchronous nature of dataflow graphs allows out-of-order execution of operators, leading to behaviors not present in the original imperative programs.
We address this challenge by developing a translation validation technique for dataflow compilers to ensure that the dataflow program has the same behavior as the original imperative program on all possible inputs and schedules of execution. We apply this method to a state-of-the-art dataflow compiler targeting the RipTide CGRA architecture. Our tool uncovers 7 compiler bugs where the compiler outputs incorrect dataflow graphs, including a data race that is otherwise hard to discover via testing. After repairing these bugs, our tool verifies the correct compilation of all programs in the RipTide benchmark suite.
1. Introduction
Even as Moore’s Law ends, many applications still demand better power efficiency, e.g., for high-performance machine learning or for ultra-low-power environmental monitoring. A flood of new hardware architectures have emerged to meet this demand, but one attractive approach that retains general-purpose programmability is a class of dataflow architectures called coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ). In a CGRA architecture, an array of processing elements (PEs) is physically laid out on the chip. Each PE can be configured to function as a high-level operator that performs an arithmetic operation, a memory operation, or a control-flow operation. These PEs are interconnected via an on-chip network, reducing communication and data movement costs compared to traditional von Neumann architectures, thus making a CGRA more power efficient.
Abstractly, CGRAs run dataflow programs, or networks of operators that execute independently and asynchronously. While CGRAs are able to run dataflow programs efficiently due to their inherent parallelism, this parallelism also introduces new, subtle correctness issues due to the potential for data races. As a result, dataflow programs are not written directly, but are instead compiled from ordinary, imperative languages. However, the gap between imperative and dataflow programming means that these compilers are complex; to maximize performance, they must use a variety of program analyses to convert imperative programs into equivalent distributed ones.
As CGRA architectures continue to emerge and rapidly evolve, we argue that formal verification — mathematically sound proofs of system correctness — should be integrated early on; past experience teaches us that belatedly retrofitting them may be costly or simply infeasible. Since CGRAs cannot execute as intended without correct dataflow programs, we begin this effort by verifying the translation from imperative programs to dataflow programs.
In this work, we propose using translation validation (pnueli-tv, ) to certify the results of dataflow compilers. In our case, translation validation amounts to checking that the output dataflow program from the compiler has the same behaviors as the input imperative program. Compared to traditional testing, translation validation provides a much stronger guarantee, as it checks that the imperative and the dataflow programs behave the same on all possible inputs. Additionally, in contrast to full compiler verification, translation validation allows the use of unverified compilation toolchains, which is crucial for the fast-developing area of CGRA architectures.
Our main challenge for leveraging translation validation is managing the asynchronicity of dataflow programs. Most translation validation works (pnueli-tv, ; necula-tv, ; kundu-tv, ; data-driven-eq, ) prove equivalence between input and output programs via (bi)simulation, or relational invariants between the states of the two programs. Since dataflow programs have exponentially more states than imperative programs (arising from asynchronous scheduling decisions), constructing this bisimulation relation directly would be quite tricky.
Instead, we utilize a two-phase approach. First, we use symbolic execution to prove that there is a bisimulation between the input imperative program and the output dataflow program on a canonical schedule. In the canonical schedule, the dataflow operators are scheduled to fire in a similar order to their counterparts in the original imperative program.
Next, we prove that any possible schedule of the dataflow program must converge to the same final state as the canonical schedule; i.e., it is confluent (term-rewriting, ). For our setting, proving confluence requires showing that the dataflow program does not contain any data races. Taking inspiration from separation logic (separation-logic, ), we augment our symbolic execution with linear permission tokens to track ownership of memory locations. Taken together with the initial bisimulation phase, we can prove that the imperative program is equivalent to the dataflow program on all possible inputs and schedules.
We have implemented this technique in WaveCert  [^1] Available on GitHub at https://github.com/secure-foundations/riptide-verification., a translation validation system targeting the state-of-the-art CGRA architecture RipTide (riptide, ). RipTide features a dataflow compiler from LLVM (llvm, ) programs to dataflow programs. Using our tool, we found 7 compiler bugs where the RipTide dataflow compiler generated incorrect dataflow programs. All of these bugs were confirmed by the developer of the RipTide compiler. One of these bugs allows the compiler to emit a dataflow program with data races, which are hard to discover via testing and costly to fix after deployment.
To summarize, our contributions in this work are:
  • A novel two-phase translation validation technique for certifying compilation results from imperative programs to dataflow programs;
  • An implementation of our technique, WaveCert, which targets the RipTide dataflow compiler; and
  • An evaluation of WaveCert on the RipTide dataflow compiler, which revealed 7 compilation bugs. Each successful verification result takes less than 30 seconds, with most taking around 10 seconds.
Examples of operators.
[…]
2. Preliminaries: CGRAs and RipTide
Dennis et al. (dennis-dataflow, ) introduced early designs for dataflow programming and architectures in the mid 1970s to allow programs to achieve a higher degree of concurrency. In the intervening years, the community has explored a wide variety of dataflow architectures (dataflow-architectures, ; yazdanpanah-dataflow, ). The desire for low-power, general-purpose programmability has recently led to a line of work on coarse-grained reconfigurable arrays (CGRAs) (cgra-survey, ).
To utilize the efficiency and parallelism in CGRAs, the predominant general-purpose way to program CGRAs is to compile from a sequential, imperative program into a dataflow graph via a dataflow compiler. This is used in CGRA projects such as PipeRench (piperench, ), WaveScalar (wavescalar, ), and more recently, RipTide (riptide, ) and Pipestitch (pipestitch, ).
In this work, we focus on the approach taken by RipTide (riptide, ), which has a dataflow compiler from arbitrary LLVM functions to dataflow programs. For an informal description of the semantics of the LLVM intermediate representation (IR), we refer the reader to the official LLVM IR manual (llvm-ir, ). In this work, we use the subset of LLVM IR that is supported by the RipTide compiler, including basic control-flow and integer arithmetic.
The dataflow program output by the RipTide compiler is a Turing-complete representation of programs. It is represented as a dataflow graph in which the nodes are called operators and edges are called channels. Semantically, channels are unbounded FIFO queues used for communication between operators. Operators can be thought of as processes running in parallel with a local state. Repeatedly, each operator: waits to dequeue (a subset of) input channels to get input values; performs a local computation, optionally changing the local state or global memory; and pushes output values to (a subset of) output channels. When an iteration of this loop is done, we say that the operator has executed or fired.
We model the execution of dataflow operators using a fully asynchronous semantics. Different CGRAs would use different scheduling strategies (cgra-survey, ), so we conservatively use asynchronicity to cover all possible schedules.
In Figure 1, we show the five most important types of operators. Reading from left to right, we first have the control flow operators, carry (C) and steer (T). The carry operator is used to support loop variables. It has two local states: in the initial state, it waits for a value from A, outputs it to O, and transitions to the loop state. In the loop state, it waits for values from B and D (decider); if D is true, it outputs O = B; otherwise it discards B and transitions to the initial state.
The steer operator is used for conditional execution. It waits for values from A and D. If D is true, it outputs O = A; otherwise, it discards the value, and outputs nothing.
To utilize the global memory, we have the load (LD) and store (ST) operators. The load operator waits for values from P (base), I (offset), and an optional signal S for memory ordering; reads the memory location P[I]; and outputs the read value to O. Similarly, the store operator waits for values from P (base), I (offset), V (value), and an optional signal S for memory ordering; stores P[I] = V in the global memory; and optionally outputs a finish signal to O.
[…]
To compile an LLVM program to a dataflow program, the RipTide compiler first maps each LLVM instruction to its corresponding dataflow operator, and then it enforces the expected control-flow and memory ordering semantics of the original program. Using a control dependency analysis, it inserts steer operators to selectively enable/disable operators depending on branch/loop conditions. Using a memory ordering analysis, it inserts dataflow dependencies between load/store operators to prevent data races.
These analyses are quite complex and use various optimizations to allow more parallelism than the original LLVM program. As shown in our evaluation in Section 6, this process can easily have bugs. Hence, in our work, we take the translation validation approach to certify the correctness of dataflow compilation.
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
1define void @test(i32* %A, i32* %B, i32 %len) {
2entry:
3  br label %header
4
5header:
6  %i = phi i32 [ 0, %entry ], [ %i_inc, %body ]
7  %cond = icmp slt i32 %i, %len
8  br i1 %cond, label %body, label %end
9
10body:
11  %idx1 = getelementptr i32, i32* %A, i32 %i
12  %A_i = load i32, i32* %idx1
13  %A_i_inc = add i32 %A_i, 1
14  %idx2 = getelementptr i32, i32* %B, i32 %i
15  store i32 %A_i_inc, i32* %idx2
16  %i_inc = add i32 %i, 1
17  br label %header
18
19end:
20  ret void
21}
(a) Input LLVM function.
(b) Output dataflow program.
An example.
Figure 2. An example of input/output programs from the RipTide dataflow compiler.
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
Cut point 2 trace.
Figure 3. A trace of execution from the dataflow cut point 2 following the canonical schedule of operators: 1, 2, 4, 7, 8, 9, 6 (which are inferred from the LLVM trace from cut point 2 to cut point 2). In addition, steer operators 3 and 5 are also fired in step 3 immediately after they are ready. At the end of the trace, the configuration matches cut point 2 again. For each configuration, we show the state of the channels (solid lines for channels with one value, and dotted lines for empty channels). The lightning symbols indicate which operators are fired in the previous step. Before the firing of an operator, the input channels of the operator should be non-empty (i.e., marked with a solid line). After firing, the input channels are emptied, and the output channels will become non-empty. The exact symbolic values in each channel are omitted, but their attached permission variables (starting with p) are marked. At the top of each configuration, we also indicate the list of permission constraints we add at each step.
[…]
Index LLVM Dataflow Correspondence
1 %entry Initial config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
2
%body
$\to$ %header
See Figure 3
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
[…]
$\bot$ %end Final config
$\mathrm{mem}_{\mathrm{llvm}}=\mathrm{mem}_{\mathrm{df}}$
Cut points for the example.
[…]
3. Overview and an Example
WaveCert performs translation validation on the RipTide dataflow compiler, which compiles LLVM programs to dataflow programs. Given the input program and the compiled output program (along with some hints generated by the compiler), WaveCert performs two checks: first, a bisimulation check, which verifies that the LLVM program is equivalent to the dataflow program on a restricted, canonical schedule of dataflow operators; and second, a confluence check, which proves that the choice of schedule for the dataflow program does not affect the final result.
If a pair of input/output programs passes both checks, the two programs are equivalent. If the bisimulation check fails, there might be a compilation bug that produced an incorrect dataflow. If the bisimulation check succeeds but the confluence check fails, there might be a data race in the output dataflow program.
Example
In Figure 1(b), we have an example input and output program from the dataflow compiler. The input LLVM function @test in Figure 1(a) contains a single loop with loop header block %header and loop body block %body. The loop increments variable %i from 0 to %len, and at each iteration, updates %B[%i] = %A[%i] + 1.
The LLVM function @test gets compiled to the dataflow program in Figure 1(b), in which most operators correspond to an LLVM instruction:
  • The comparison operator 4 corresponds to line 7;
  • The load operator 7 corresponds to line 12;
  • The add operator 8 corresponds to line 13;
  • The store operator 9 corresponds to line 15; and
  • The add operator 6 corresponds to line 16.
The compiler also inserts additional operators to faithfully implement the sequential semantics of the LLVM program. The steer operators marked with T enforce that the operators corresponding to the loop body only execute when the loop condition is true. The carry operator 1 is used for the loop induction variable %i (line 6). The carry operator 2 is used for a loop variable inserted by the compiler to enforce a memory dependency: since arrays A and B could overlap, we have an additional dataflow path from the store operator, passing through the carry operator 2 and the steer operator 5, to the load operator, in order to enforce that the load in the next iteration waits until the previous store has finished.
Bisimulation Check
The first step in WaveCert’s translation validation procedure is to check that the LLVM function is bisimilar to the dataflow program on a canonical schedule of operators, where the operators are executed in the same order as their LLVM instruction counterparts.
To check a bisimulation, we need to construct a relation between the states of the LLVM and dataflow programs. WaveCert constructs this relation by placing corresponding cut points on both the LLVM and the dataflow sides. A cut point symbolically describes a set of pairs of LLVM and dataflow configurations satisfying a particular correspondence condition, and a list of cut points together describes the bisimulation relation. Figure 4 shows the list of cut points for the example. WaveCert places cut point 1 for the initial configurations, cut point 2 for the loop structure, and a final cut point $\bot$ for all final configurations. For each pair of LLVM and dataflow cut points, WaveCert also infers a list of equations expressing the correspondence of symbolic variables in these cut points.
To automatically check that the proposed relation in Figure 4 is indeed a bisimulation, we perform symbolic execution from all pairs of cut points (except for $\bot$) until they reach another pair of cut points. Symbolic execution (symbolic-execution, ) is similar to concrete execution of programs, except with concrete values replaced by symbolic expressions. In this way, we can execute all possible concrete instances of a cut point altogether. On the LLVM side, we first symbolically execute cut point 1, which turns into two symbolic branches: reaching %header after one loop iteration (cut point 2), or failing the loop condition and reaching %end (cut point $\bot$). For LLVM cut point 2, we also have two branches: one reaching cut point 2 again after a loop iteration, the other failing the loop condition and reaching cut point $\bot$.
For the dataflow side, we have a correspondence between a subset of LLVM instructions and a subset of dataflow operators, which is automatically generated by the dataflow compiler. We use this mapping to infer the order in which we fire the dataflow operators, i.e., the canonical schedule. For example, from LLVM cut point 2, there are two branches going to cut point 2 and cut point $\bot$. These two branches have two traces of LLVM instructions (excluding br):
  • For the first branch to cut point 2: lines 6, 7, 11 - 16;
  • For the second branch to cut point $\bot$: lines 6, 7.
For the first branch, the list of LLVM instructions map to dataflow operators 1, 2, 4, 7, 8, 9, 6 (excluding ones that do not have a corresponding operator; operator 2 corresponds to a PHI instruction implicitly inserted during compilation for memory ordering). LLVM instructions in the second branch map to dataflow operators 1, 2, and 4. Therefore, for dataflow cut point 2, we execute these two traces of dataflow operators (and also any steer operator that can be executed), which gives us two dataflow configurations. Figure 3 shows the first branch from cut point 2 to itself. At the end of the execution, we can see that the configuration matches the cut point 2 configuration we started with, except with potentially different memory configurations and symbolic variables (e.g. i + 1 vs. i in the channel from operator 6 to 1).
[…]
$\mathrm{mem}_{\mathrm{llvm}}[\texttt{\%B}[\texttt{\%i}]\mapsto\texttt{\%A}[\texttt{\%i}]+1]=\mathrm{mem}_{\mathrm{df}}[\texttt{B}[\texttt{i}]\mapsto\texttt{A}[\texttt{i}]+1]$
%i + 1=i + 1%i + $=\texttt{i + 1}$ i + 1
If this check succeeds for all dataflow and LLVM cut points, we soundly conclude that we have obtained a bisimulation between the LLVM program and the dataflow program (under the deterministic canonical schedule).
Confluence Check
Once we have established a bisimulation between the LLVM function and the canonical schedule of the dataflow program, we check that the dataflow program is confluent: all possible schedules converge to the canonical schedule in the final state (or synchronize with it infinitely often if the program is non-terminating).
[…]
We attach a permission variable (representing a permission token that we do not know a priori) to each value in the channels of dataflow configurations in each symbolic branch. Figure 3 shows an example of this attachment of permission variables and permission constraints. At the starting state cut point 2, we attach free permission variables $p_{1},p_{2},p_{3},p_{4}$ to values in the configuration. After the first step in which operators 1 and 2 fire, these permission variables are consumed by the two carry operators and we generate fresh permission variables $p_{5},p_{6},p_{7}$ for the new output values. We require that the permission variables are used in a linear fashion: operators cannot duplicate or generate permission tokens, instead, they have to be obtained from the inputs. So for the first step in Figure 3, we add two constraints $p_{5}+p_{6}\leq p_{1}+p_{3}$ and $p_{7}\leq p_{2}+p_{4}$ to say that the output permissions have be contained in the input permissions. Furthermore, if the operator is a load or a store, such as in Figure 3 steps 4 and 6, we require that the suitable permission (read for load, and write for store) should be present in the input permissions (e.g. $\texttt{write}\leq p_{13}+p_{17}$ at step 6 in Figure 3).
Finally, when the execution is finished and has either hit a final state or another cut point, we have collected a list of constraints involving a set of permission variables. If the configuration is matched to a cut point, we add additional constraints to make sure that the assignment of tokens at the cut point is consistent. For example in Figure 3, for the last configuration, we would add these additional constraints: $p_{1}=p_{19}$, $p_{2}=p_{18}$, $p_{3}=p_{8}$, $p_{4}=p_{9}$. We then solve for the satisfiability of these constraints with respect to the permission algebra ($(\{0,\texttt{read},\texttt{write}\},+,\leq)$ we have defined above.
If there is a satisfying assignment of the permission variables with permission tokens, then intuitively, the operators in the dataflow program are consistent in memory accesses, and there will not be any data races. In Section 5, we show that this implies that any possible schedule will converge to the canonical schedule we use. For the example in Figure 3, a satisfying assignment is $p_{2}=p_{7}=p_{15}=p_{16}=p_{17}=p_{18}=\texttt{write}$ with all other variables set to 00. This assignment essentially passes the write permission across the loop of operators 2, 5, 7, 8, 9.
4. Bisimulation on the Canonical Schedule
Now, we detail the first half of our translation validation technique, which proves via cut-bisimulation that the LLVM program is equivalent to the dataflow program when restricted to a canonical schedule.
Given a dataflow program and LLVM program, their semantics are defined by transition systems $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ and $(\mathcal{I}_{\mathrm{llvm}},\mathcal{C}_{\mathrm{llvm}},\to_{\mathrm{llvm}})$, respectively, where $\mathcal{I}_{\mathrm{df}}$ and $\mathcal{I}_{\mathrm{llvm}}$ are the sets of initial configurations and $\mathcal{C}_{\mathrm{df}}$ and $\mathcal{C}_{\mathrm{llvm}}$ are the sets of all configurations.
We use $\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})$ to denote the state of the memory in a dataflow configuration $c_{\mathrm{df}}$, and similarly $\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})$ for the memory in an LLVM configuration $c_{\mathrm{llvm}}$. We will omit the subscripts when it is clear from the context. For the dataflow semantics, we use $\to^{o}_{\mathrm{df}}$ to denote the (deterministic) firing of a particular operator $o$. For a transition relation $\to$, we use $\to^{*}$ for the reflexive-transitive closure of $\to$, and $\to^{+}$ for the transitive closure of $\to$. A configuration $c$ is terminating/final if there exists no configuration $c^{\prime}$ such that $c\to c^{\prime}$.
One important tool that we use is symbolic execution, which is similar to concrete execution but performed with symbolic expressions in the configuration. In the following, we establish some notation for symbolic execution. Let $\mathbf{x}$ be a list of variables $\mathbf{x}=(x_{1},\dots,x_{n})$. Let $V$ be the set of all bit-vectors of any lengths. We use $V(\mathbf{x})$ to denote the set of symbolic expressions constructed from constants in $V$, variables in $\mathbf{x}$, and any bit-vector operation we could perform in our semantics (e.g. 32-bit zero bit-vector $\texttt{0}_{32}$, or terms like $x_{1}+\texttt{1}_{32}$ are in $V(\mathbf{x})$). For $s\in\{\mathrm{df},\mathrm{llvm}\}$, we use $\mathcal{C}_{s}(\mathbf{x})$ to represent the set of symbolic configurations with free variables $\mathbf{x}$ in the semantics $s$. A symbolic configuration $c(\mathbf{x})\in\mathcal{C}_{s}(\mathbf{x})$ is similar to a configuration but over symbolic expressions $V(\mathbf{x})$ instead of only $V$ (for example, a symbolic dataflow configuration can have symbolic expressions such as $x_{1}+\texttt{1}_{32}$ in the channels; a symbolic LLVM configuration can have symbolic expressions stored in variables as opposed to only concrete values in $V$). If we have a substitution $\sigma:\{x_{1},\dots,x_{n}\}\to V(\mathbf{y})$ from $\mathbf{x}$ to symbolic expressions over some other list of variables $\mathbf{y}$, we denote $c(\sigma)\in\mathcal{C}_{s}(\mathbf{y})$ to be the result of the substitution. We also lift the transition relation $\to_{s}$ (for $s\in\{\mathrm{df},\mathrm{llvm}\}$) to symbolic configurations: we use $c(\mathbf{x})\to_{s}c^{\prime}(\mathbf{x})\mid\varphi(\mathbf{x})$ to denote the symbolic execution from $c(\mathbf{x})$ to $c^{\prime}(\mathbf{x})$ with path condition $\varphi(\mathbf{x})$.
The dataflow transition relation $\to_{\mathrm{df}}$ could be nondeterministic (a configuration can have multiple outgoing transitions), while $\to_{\mathrm{llvm}}$ is deterministic. However, for this section, we only care about a canonical schedule for the execution of operators. We denote the transition relation restricted to this canonical schedule by $\to_{\mathrm{df\mid canon}}\subseteq\to_{\mathrm{df}}$, and we focus on this deterministic transition for this section. The exact meaning of this canonical schedule will be explained in Section 4.2, but essentially, it executes dataflow operators in the same order their corresponding LLVM instructions would execute.
The equivalence property that we want to check is that for any pair of concrete initial states $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$,
  • $c_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for a final state $c^{\prime}_{\mathrm{df}}$ iff $c_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for a final state $c^{\prime}_{\mathrm{llvm}}$, in which case $\mathrm{mem}(c^{\prime}_{\mathrm{df}})=\mathrm{mem}(c^{\prime}_{\mathrm{llvm}})$.
  • $c_{\mathrm{df}}$ is non-terminating iff $c_{\mathrm{llvm}}$ is non-terminating, and there are infinitely many pairs of states in their traces with equal memory.
Since the dataflow and LLVM traces could both be unbounded, naive symbolic execution from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$ may not terminate. So we need to prove a coinductive invariant about the pairs of traces starting from $(\mathcal{I}_{\mathrm{df}},\mathcal{I}_{\mathrm{llvm}})$, and such an invariant is a cut-bisimulation relation.
Definition 0 (Cut-bisimulation).
Let $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ be a binary relation. We say that $\mathcal{R}$ is a cut-bisimulation iff for any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$, if $c_{\mathrm{df}}\to_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$ for some $c^{\prime}_{\mathrm{df}}$, then there are $(c^{\prime\prime}_{\mathrm{df}},c^{\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{df}}\to^{*}_{\mathrm{df\mid canon}}c^{\prime\prime}_{\mathrm{df}}$ and $c_{\mathrm{llvm}}\to^{+}_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$. Conversely, if $c_{\mathrm{llvm}}\to_{\mathrm{llvm}}c^{\prime}_{\mathrm{llvm}}$ for some $c^{\prime}_{\mathrm{llvm}}$, then there are $(c^{\prime}_{\mathrm{df}},c^{\prime\prime}_{\mathrm{llvm}})\in\mathcal{R}$ such that $c^{\prime}_{\mathrm{llvm}}\to^{*}_{\mathrm{llvm}}c^{\prime\prime}_{\mathrm{llvm}}$ and $c_{\mathrm{df}}\to^{+}_{\mathrm{df\mid canon}}c^{\prime}_{\mathrm{df}}$.
This is a weaker variant of the usual notion of bisimulation where we only allow one step before synchronizing instead of allowing one or more steps. Work by Kasampalis et al. (keq, ) provides more detail about this variant of bisimulation.
To check equivalence, we look for cut-bisimulations that are memory-synchronizing:
Definition 0 ().
A binary relation $\mathcal{R}\subseteq\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}$ is memory-synchronizing iff
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\subset\mathcal{R}$, $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}$ with $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
  • For any $(c_{\mathrm{df}},c_{\mathrm{llvm}})$ with $c_{\mathrm{df}},c_{\mathrm{llvm}}$ final and $\mathrm{mem}(c_{\mathrm{df}})=\mathrm{mem}(c_{\mathrm{llvm}})$, $(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{R}$.
If there exists a cut-bisimulation $\mathcal{R}$ that is memory synchronizing, then we can conclude via a coinductive argument that the two programs are equivalent.
In the following subsections, we first describe how we represent a cut-bisimulation relation via cut points (Section 4.1), and how to automatically check that a proposed relation is actually a cut-bisimulation via symbolic execution (Section 4.2). Then in Section 4.3, we describe how the cut points are inferred using hints from the compiler.
4.1. Describing a Cut-Bisimulation via Cut Points
To finitely describe a proposed cut-bisimulation, we use a list of pairs of dataflow and LLVM cut points. A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and LLVM configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$. In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equations between variables in $\mathbf{x}$ and $\mathbf{y}$. In particular, we require $\mathrm{mem}(c_{\mathrm{df}}(\mathbf{x}))=\mathrm{mem}(c_{\mathrm{llvm}}(\mathbf{y}))$ to be in $\varphi(\mathbf{x},\mathbf{y})$ for any cut point. Semantically, each cut point describes the relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$. For a list of $n$ cut points
$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
the relation they describe is the union of the subrelation each describes: $\mathcal{R}(P^{1},\dots,P^{n}):=\bigcup^{n}_{i=1}\mathcal{R}(P^{i})$.
To ensure that the cut-bisimulation we propose satisfies the property in Definition 2, we also require that the first cut point must be exactly the pairs of initial configurations with equal memory. That is $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}\mid\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Furthermore, we require that the last cut point $P^{n}$ must be the pairs of final configurations with equal memory, that is $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid c_{\mathrm{df}},c_{\mathrm{llvm}}\text{ final},\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Then $\mathcal{R}(P^{1},\dots,P^{n})$ satisfies Definition 2.
4.2. Checking Cut-Bisimulation
Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 2, we need to check that it is indeed a cut-bisimulation.
To achieve this, we perform symbolic execution from each cut point and try to reach another cut point. Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$. We perform symbolic execution from $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ and $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ until all of their symbolic branches hit another cut point. For the LLVM side, we get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,1}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},1}(\mathbf{y^{i}})$
$\dots$
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,k_{i}}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},k_{i}}(\mathbf{y^{i}})$
where each right hand side $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{llvm}}(\sigma)$ of some other LLVM cut point $c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$. Furthermore, since our LLVM semantics is deterministic, the path conditions should be disjoint, i.e., for any $l_{1},l_{2}\in\{1,\dots,k_{i}\}$, if $l_{1}\neq l_{2}$, $\psi_{\mathrm{llvm},l_{1}}(\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l_{2}}(\mathbf{y^{i}})$ is unsatisfiable.
For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, we try to replicate the execution of LLVM branches on the dataflow side. For each LLVM branch $l\in\{1,\dots,k_{i}\}$, we can get a trace of LLVM instructions that are executed from $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ to $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$. Using hints from the compiler, we know that a subset of these LLVM instructions map to a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$. For each LLVM branch, to produce the corresponding dataflow branch, we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$:
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}*}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}*}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$
where each $\to^{o*}_{\mathrm{df}}$ denotes the dataflow transition by first firing the operator $o$, and then firing any available steer operator until none is fireable anymore. The canonical schedule $\to_{\mathrm{df\mid canon}}$ is defined by this order of execution of operators.
Therefore, by replicating the execution of the LLVM branches from cut point $i$, we get a list of dataflow branches:
(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
$\dots$
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
where each path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$. We check that the dataflow symbolic execution from cut point $i$ has covered all concrete cases, i.e., all possible cases of $\mathbf{x^{i}}$, by checking the unsatisfiability of $\neg\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$.
Finally, we check that each pair of dataflow and LLVM branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))$ matches some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
$\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to V(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{llvm}}:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$ such that the following conditions hold:
  • The dataflow branch is an instance of the target dataflow cut point: $c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})=c^{j}_{\mathrm{df}}(\sigma^{i,l}_{\mathrm{df}})$.
  • The LLVM branch is an instance of the target LLVM cut point: $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})=c^{j}_{\mathrm{llvm}}(\sigma^{i,l}_{\mathrm{llvm}})$.
  • The correspondence condition at the target cut point is satisfied given the source cut point correspondence and the path conditions:
    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{llvm}})$
If the check described in this section is passed for all cut points $i$ and all branches $1,\dots,k_{i}$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-bisimulation.
4.3. Heuristics to Infer Cut Points
The final step is to actually infer the LLVM and dataflow cut points. In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable (rogers-87, ), so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
The placement of cut points is important for the symbolic execution to terminate. It is in theory correct (for terminating dataflow and LLVM programs) to have exactly two cut points: one at the initial configurations, and one for all final configurations. But the symbolic execution will try to simulate all possible traces together, which are unbounded, and the check may not terminate. To ensure termination, on the LLVM side, we place a cut point at each of the following program points:
  • The entry of the function (i.e. the initial configuration);
  • The back edge of each loop header block; and
  • The exit program point.
For the dataflow program, we need to place the first cut point at the initial configuration of the dataflow program, and the last cut point at the final configuration. For loops, ideally, we want to place the cut points at the corresponding “program points” at loop headers. However, due to the lack of control-flow in the dataflow program, the form of the corresponding dataflow configurations at these “program points” is hard to infer statically. Instead of generating the dataflow cut points ahead of time, we dynamically infer them during symbolic execution of dataflow branches. For example, after mirroring the LLVM branch execution in a dataflow branch such as in Equation 1, if the dataflow cut point corresponding to the LLVM cut point has not been inferred yet, we generalize the symbolic configuration on the RHS of the dataflow branch Equation 1 to a dataflow cut point by replacing symbolic expressions with fresh variables. The correspondence condition is then inferred using compiler hints to match fresh variables in the new dataflow cut point to an LLVM variable defined by some LLVM instruction.
Note that we do not need these heuristics to be sound. As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-bisimulation check in Section 4.2 succeeds, the soundness of the cut-bisimulation follows. While we do not prove these heuristics correct, they work well in practice and cover all test cases in our evaluation (Section 6).
5. Confluence Checking using Linear Permission Tokens
After checking cut-bisimilarity between the LLVM program and the dataflow program on the canonical schedule, we still need to prove that this canonical schedule is general. Indeed, since the dataflow program is asynchronous and nondeterministic, other choices of schedule may lead to data races and inconsistent final states.
Therefore, in addition to finding a bisimulation on the canonical schedule, we want to prove a confluence property on the dataflow program:
Definition 0 ().
We say $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is (ground) confluent iff for any initial configuration $c\in\mathcal{I}_{\mathrm{df}}$, if $c\to^{*}_{\mathrm{df}}c_{1}$ and $c\to^{*}_{\mathrm{df}}c_{2}$, then there is $c^{\prime}\in\mathcal{C}_{\mathrm{df}}$ with $c_{1}\to^{*}_{\mathrm{df}}c^{\prime}$ and $c_{2}\to^{*}_{\mathrm{df}}c^{\prime}$.
In this section, we present a technique using what we call linear permission tokens to ensure that the dataflow program is confluent. Linear permission tokens are, in essence, a dynamic variant of the fractional permissions studied in the concurrent separation logic (fractional-perm, ; concurrent-separation-logic, ). In our case, their linear usage is dynamically enforced and the exact placement of permission tokens is resolved after symbolic execution.
We associate each value present in a dataflow execution trace with a permission token in the form of $\texttt{write}\ l$, $\texttt{read}\ l$, or any formal sum of them, for any memory location $l$. These tokens are used by dataflow operators in a linear fashion: they cannot be replicated, and they are disjoint in the initial state. A write token can be split into $n$ many read tokens (for a predetermined parameter $n$), and inversely, $n$ of such read tokens can be combined to get back a write token. As shown later in this section, if there is a consistent assignment of these linear permission tokens, we can conclude that there are no data races possible in any schedule, and thus all final states have to be the same.
In Section 5.1, we prove that the technique works in a bounded setting for a concrete trace $\tau$: if $\tau$ has a consistent assignment of permission tokens, then for any other trace $\tau^{\prime}$ bounded by $\tau$ (meaning that for each operator, $\tau^{\prime}$ executes it for an equal or smaller number of times than $\tau$), $\tau^{\prime}$ will converge to the final configuration of $\tau$.
Then in Section 5.2, we explain how this bounded confluence check can be extended to a full confluence check for the entire dataflow program by using the cut points that we have placed for the dataflow program.
5.1. Confluence on Bounded Traces
In this section, we show that if we have a $n$-step trace $c_{1}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}$ (where $o_{1},\dots,o_{n}$ are the operators fired in each step) with a consistent assignment of permission tokens to each value, then for any other trace $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ (not necessarily with a consistent assignment of permission tokens) such that the multiset $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, we have $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
Let us first describe the permission tokens and their assignment to values. The permission tokens are drawn from the permission algebra defined below.
Definition 0 (Permission Algebra).
Let $L$ be a set of memory locations and $k$ be a positive integer. The permission algebra is a partial algebraic structure $(P(L,k),0,+,\leq)$ where
  • $(P(L,k),0,+)$ is the term algebra of the formal sums $\texttt{read}\ l_{1}+\dots+\texttt{read}\ l_{n}$, with $l_{1},\dots,l_{n}\in L$, modulo the commutativity and associativity of $+$, and we require that for a location $l$, at most $k$ copies of $\texttt{read}\ l$ is present in the formal sum. We call $p+q$ the disjoint union of $p,q\in P(L,k)$. We also denote
    $\texttt{write}\ l:=\underbrace{\texttt{read}\ l+\dots+\texttt{read}\ l}_{k\text{ copies}}$
  • $P(L,k)$ is partially ordered by $\leq$ with $p\leq q$ (reads $q$ contains $p$) iff for any $l$, the number of occurrences of $\texttt{read}\ l$ in $p$ is less than or equal to that in $q$.
In Boyland’s terminology (fractional-perm, ), $P(L,k)$ is a finite token model of size $k$ of fractional permissions. For example, in $P(\{\texttt{A},\texttt{B}\},2)$, we have that
[…]
Definition 0 ().
We say that the permissions $p_{1},\dots,p_{n}\in P(L,k)$ are disjoint iff for any $i\neq j$, $p_{i}+p_{j}$ is defined.
Now let us define an assignment of permission tokens to values in a configuration. Let $E$ denote the set of channel names. For a configuration $c\in\mathcal{C}_{\mathrm{df}}$ and a channel $e\in E$, let $\mathrm{Channel}(c,e)$ denote the state of the channel $e$ in $c$, which is a string over $V$ representing the values in the channel $e$.
Definition 0 (Permission Augmentation).
Let $c\in\mathcal{C}_{\mathrm{df}}$ be a dataflow configuration. A permission augmentation of $c$ is a partial map $t:E\times\mathbb{N}\to P(L,k)$ such that for any $e\in E$, $t(e,n)$ is defined iff $n\in\{1,\dots,|\mathrm{Channel}(e)|\}$. We call $(c,t)$ a permission-augmented configuration.
Intuitively, $t(e,n)$ is the permission we attach to the $n$-th value in the channel $e$. In the following, we define when such a permission augmentation is consistent (i.e., the use of permissions is linear).
Definition 0 (Consistent Augmentation).
We call $(c,t)$ a consistent (permission-augmented) configuration if the permissions in the image of $t$ are disjoint.
Definition 0 (Consistent Transition).
Let $c\to^{o}_{\mathrm{df}}c^{\prime}$ be a transition. Let $t,t^{\prime}$ be two permission augmentations for $c,c^{\prime}$, respectively. Let $p_{1},\dots,p_{n}$ be the permissions in $t$ attached to input values of $o$ and $q_{1},\dots,q_{m}$ be the permissions in $t^{\prime}$ attached to output values of $o$. We say $(c,t)\to^{o}_{\mathrm{df}}(c^{\prime},t^{\prime})$ is a consistent (permission-augmented) transition iff:
  • The $t$ and $t^{\prime}$ have the same assignment of permissions except for the channels changed by the operator $o$.
  • (Linearity) $q_{1}+\dots+q_{m}\leq p_{1}+\dots+p_{n}$.
  • (Load Permission) If $o$ is a load operator on a memory location $l\in L$, then $\texttt{read}\ l\leq p_{1}+\dots+p_{n}$.
  • (Store Permission) If $o$ is a store operator on a memory location $l\in L$, then $\texttt{write}\ l\leq p_{1}+\dots+p_{n}$.
Definition 0 (Consistent Trace).
Let $c_{1}\to_{\mathrm{df}}\dots\to_{\mathrm{df}}c_{n}$ be a trace. Let $t_{1},\dots,t_{n}$ be permission augmentations to $c_{1},\dots,c_{n}$, respectively. We say that $(c_{1},t_{1})\to_{\mathrm{df}}\dots\to_{\mathrm{df}}(c_{n},t_{n})$ is a consistent (permission-augmented) trace iff:
  • $(c_{1},t_{1})$ is consistent
  • $(c_{i},t_{i})\to_{\mathrm{df}}(c_{i+1},t_{i+1})$ is consistent for $i\in\{1,\dots,n\}$.
The following lemma allows us to swap the order of execution of operators in two consistent steps.
Lemma 8 ().
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}(c_{2},t_{2})\to^{o_{2}}_{\mathrm{df}}(c_{3},t_{3})$ be a consistent trace. Suppose $c_{1}\to^{o_{2}}_{\mathrm{df}}c^{\prime}_{2}$ for some configuration $c^{\prime}_{2}$. Then $c^{\prime}_{2}\to^{o_{1}}_{\mathrm{df}}c_{3}$ and there is a permission augmentation $t^{\prime}_{2}$ such that $(c_{1},t_{1})\to^{o_{2}}_{\mathrm{df}}(c^{\prime}_{2},t^{\prime}_{2})\to^{o_{1}}_{\mathrm{df}}(c_{3},t_{3})$ is consistent.
Proof Sketech.
If one of $o_{1}$ and $o_{2}$ is neither load nor store, or if both of them are loads, the result will always be consistent.
If $o_{1}$ and $o_{2}$ are both stores, and they are the same node, then it must be the case that $c^{\prime}_{2}=c_{2}$ and we can use $t^{\prime}_{2}=t_{2}$.
Otherwise, assume both $o_{1}$ and $o_{2}$ are load/store operators, and they are different nodes in the dataflow graph. Since they can both fire at $c_{1}$, and they are different nodes, their input values must be attached with disjoint permissions. Therefore, $o_{1}$ and $o_{2}$ have no dependencies, and they will give the same result no matter in which order we execute them. ∎
Inductively applying this lemma, we can swap the execution of an operator all the way to the beginning if it can fire and has not fired yet.
Lemma 9 ().
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}(c_{n+1},t_{n+1})$ be a consistent trace. Suppose $o_{n}$ is not in the set $\{o_{1},\dots,o_{n-1}\}$ and $c_{1}\to^{o_{n}}_{\mathrm{df}}c^{\prime}_{2}$ for some $c^{\prime}_{2}$, then $c_{1}\to^{o_{n}}_{\mathrm{df}}c^{\prime}_{2}\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n-1}}_{\mathrm{df}}c_{k}$ for some $c^{\prime}_{2}$ and there is a consistent augmentation for this trace.
Applying the lemma above inductively, we have the following.
Theorem 10 (Bounded Confluence for Consistent Traces).
Let $(c_{1},t_{1})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}(c_{n+1},t_{n+1})$ be a consistent trace. Let $c_{1}\to^{o^{\prime}_{1}}_{\mathrm{df}}\dots\to^{o^{\prime}_{m}}_{\mathrm{df}}c_{m+1}$ be another trace. If we have the multiset inclusion $\{o^{\prime}_{1},\dots,o^{\prime}_{m}\}\subseteq\{o_{1},\dots,o_{n}\}$, then $c_{m+1}\to^{*}_{\mathrm{df}}c_{n+1}$.
Practically, with this bounded confluence theorem, we can do bounded model checking for the confluence property: we perform symbolic execution to obtain a trace of symbolic configurations $c_{1}(\mathbf{x})\to^{o_{1}}_{\mathrm{df}}\dots\to^{o_{n}}_{\mathrm{df}}c_{n+1}(\mathbf{x})$, and then we can solve for a consistent augmentation $(t_{1},\dots,t_{n+1})$ for this trace using an SMT (z3, ) solver. If there is a solution, then it follows from Theorem 10 that any other trace (firing operators only in the multiset $\{o_{1},\dots,o_{n}\}$, counting multiplicity) would still converge to $c_{n+1}(\mathbf{x})$ in the end.
5.2. Full Confluence Checking using Cut Points
To extend the bounded confluence results in Section 5.1 to checking confluence for the entire dataflow program with unbounded traces, we need to finitely describe and check a consistent pattern of permission augmentations for all traces. For this purpose, we use the cut points selected in Section 4.
Let $P^{1}_{\mathrm{df}}(\mathbf{x^{1}}),\dots,P^{n}_{\mathrm{df}}(\mathbf{x^{n}})$ be a list of cut points. During the bisimulation check in Section 4.2, we have verified that starting from any cut point $P^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, any branch of symbolic execution reaches an instance of another cut point $P^{j}_{\mathrm{df}}(\mathbf{x^{j}})$. A consequence of this check is that all the symbolic branches from all cut points “cover” any complete trace (traces from an initial configuration to a final configuration). In other words, any complete concrete trace from the initial configuration is the concatenation of instances of symbolic branches.
Thus, if we are able to find a consistent permission augmentation for the trace of each symbolic branch during the bisimulation check, then any concrete trace in $\to_{\mathrm{df\mid canon}}$ would have a consistent permission augmentation, which implies that any other schedule will converge to $\to_{\mathrm{df\mid canon}}$, and the confluence of the entire dataflow program follows.
Theorem 11 ().
If all symbolic dataflow branches produced in Section 4.2 have consistent permission augmentations, then $(\mathcal{I}_{\mathrm{df}},\mathcal{C}_{\mathrm{df}},\to_{\mathrm{df}})$ is confluent.
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
Name LOC #OP #PC Bisim. (sec) Conf. (sec)
bfs 48 48 3634 0.45 7.79
dconv 24 65 8134 0.96 11.87
dfs 45 49 3762 0.46 8.51
dmm 16 56 4220 0.71 6.31
dmv 10 31 1234 0.22 1.69
fft 29 70 2690 0.45 6.04
sconv 22 56 4272 0.71 10.89
smm 22 49 3967 0.53 12.08
smv 12 25 1103 0.19 2.36
sort 23 57 6554 $\times$ 0.64 3.62
nn_conv 43 61 12154 0.78 15.65
nn_fc 24 36 2632 0.31 3.43
nn_norm 15 19 316 0.07 0.17
nn_pool 33 52 8165 0.74 7.11
nn_relu 16 19 844 0.11 0.44
nn_vadd 9 15 244 0.06 0.15
Dither 21 31 2553 0.32 2.36
SpMSpMd 42 60 7693 0.93 28.45
SpMSpVd 48 40 3182 0.4 10.22
SpSlice 33 31 2616 $\times$ 0.34 7.4
Evaluation results.
Figure 5. Evaluation results on a benchmark of 20 programs. From left to right, the columns are names of the test cases (Name), lines of code in the original C program (LOC), numbers of dataflow operators (#OP), numbers of permission constraints (#PC), time spent for bisimulation check in seconds (Bisim.), time spent for confluence check in seconds (Conf.). All test cases are written originally in C and then compiled to LLVM via Clang (clang, ). The first 10 programs are benchmark programs from the RipTide work. The nn_* tests implement neural network inference. The last four are selected from a follow-up work to RipTide called Pipestitch (pipestitch, ). In the case of sort and SpSlice, the bisimulation check fails, because they contain instances of the same compiler bug. In all other cases, the bisimulation and confluence checks succeed.
6. Implementation and Evaluation
We have implemented our translation validation technique in the tool WaveCert  [^2] Available on GitHub at https://github.com/secure-foundations/riptide-verification.. We first implemented symbolic executors for (a subset of) LLVM and dataflow programs in Python (z3, ). Based on these, we implemented the bisimulation check in Section 4 and the confluence check in Section 5.
WaveCert uses the Z3 SMT solver (z3, ) to discharge the verification conditions it generates. In the bisimulation check, WaveCert encodes the validity of the correspondence conditions at each cut point into Z3 queries. In the confluence check, WaveCert produces a list of permission constraints and checks if these constraints are satisfiable with respect to the finite permission algebra (Definition 2) by encoding each permission variable as a set of Boolean variables and the list of permission constraints as a SAT query.
In order to evaluate WaveCert’s capability to certify compilation to real dataflow programs, we have applied WaveCert to a benchmark of 20 programs consisting of the benchmark programs in RipTide (riptide, ) and Pipestitch (pipestitch, ), as well as some programs implementing neural network inference. Figure 5 shows some statistics about the benchmark programs and how WaveCert performs. All tests are performed on a laptop with an M1 Max processor and 64 GiB of RAM.
In all of the benchmark programs, the bisimulation check either succeeds or finds an actual bug (such as in sort and SpSlice, which we discuss further in Section 6.1). This shows that our heuristics for inferring dataflow cut points works well on the canonical schedule. Our bisimulation check generates verification conditions in quantifier-free theories of arrays and integers without non-linear arithmetic, so all bisimulation checks finish quickly within a second.
The confluence check succeeds in all examples. For sort, we manually fixed the incorrect dataflow program produced by the compiler, after which the confluence check passes.
Figure 6. Examples of two compiler bugs.
Examples of two compiler bugs.
Figure 6. Examples of two compiler bugs.
6.1. Compiler Bugs
During the testing of WaveCert, we found 7 bugs in the RipTide compiler. Three of these bugs were discovered on the test cases sort, SpSplice, and their variations; and the rest four bugs were found on smaller test cases that we manually constructed. A developer of the RipTide compiler has confirmed these issues.
In the following, we discuss two examples of these bugs to show that the bisimulation check and the confluence check are both effective in catching compilation bugs.
Incorrect signed extension
WaveCert found a compilation bug involving integer signed vs. unsigned extension in the sort test case. A simplified LLVM program used to reproduce this issue is shown in the top left of Figure 6, and the output dataflow program is shown in the top right of Figure 6. The issue itself is straightforward: in the LLVM code, a 1-bit xor is done with a 1-bit constant 1, and then the result is zero-extended to a 32-bit integer. On the dataflow side, the dataflow compiler tries to convert all operations into 32-bits, the word width supported on the RipTide architecture. However, during the process, the compiler assumes that all constants should be sign-extended. As a result, it extends the 1-bit 1 to a 32-bit -1, which is incorrect in this case. Our bisimulation check catches this bug when verifying the correspondence condition at the final cut point, where the memory configurations do not match.
Incorrect memory ordering with potential data races
Our confluence check catches a compilation issue when compiling from the C program shown in the bottom left of Figure 6 to the dataflow program in the bottom right of Figure 6. In the C program, we have a nested loop. In each iteration of the outer loop, we first set A[i] = i, and then in the inner loop that follows, we copy A[i] into A[i + 1], …, A[n - 1]. In the output dataflow program, the compiler allows the load and store operators in the inner loop to pipeline using the invariant operator (marked I) to repeatedly send the signal to enable load. However, this causes a data race, because in the inner loop, a load in the second iteration accesses the same memory location as the store in the first iteration.
Our confluence check fails as the list of permission constraints we collected are unsatisfiable. WaveCert also outputs an unsat core of 57 constraints, a subset of the 420 constraints generated, that causes the unsatisfiability. In particular, the unsat core includes two ordering constraints: read  A is contained in the disjoint union of input tokens of the inner loop load and write  A is contained in the disjoint union of input tokens of the inner loop store. This indicates that the main conflict stems from the inner loop load and store.
7. Limitations and Future Work
Our translation validation technique still has some limitations, and we plan to address them in future work.
Optimized dataflow graphs
Our technique may fail on more optimized programs. The RipTide compiler employs some optimizations currently not supported by WaveCert:
  • Deduplication of operators, which combines arithmetic operators with the same inputs.
  • Streamification, which uses a single stream operator to generate a loop induction variable (as opposed to using, e.g., carry, comparison, and addition operators).
  • Array dependency analysis, which removes unnecessary orderings between loads and stores across iterations that access different indices of an array.
The first two optimizations are not supported by our bisimulation check because we require a one-to-one correspondence between LLVM instructions and dataflow operators for the canonical schedule. In future work, we can relax this requirement and allow multiple LLVM instructions to map to one dataflow operator, and we then only fire the operator if all of the corresponding LLVM instructions have fired.
The third optimization may lead to a dataflow program failing our confluence check. To solve this, we need finer granularity in permission tokens. Currently, the read and write permissions are for the entire memory region, such as an array A. To allow memory operations on A[i] and A[i + 1] to run in parallel, where i is the loop induction variable, we need to re-formulate the permission algebra to allow slices of a permission token on a specific range of indices. This also requires a form of “dependent” permission tokens that can depend on the values in the channels.
Liveness properties
Our semantics of the dataflow program uses unbounded FIFO queues to model dataflow channels. In this model, we do not need to consider liveness issues such as deadlocks, because an operator can always fire once the input values are ready. However, this is not true for the actual hardware, which has a limit on the queue size.
For future work, we plan to explore model checking (model-checking, ) techniques to check for deadlocks.
Lack of architecture-specific details
This work targets the compilation from imperative programs to an abstract version of dataflow programs. While our technique can ensure the correctness of this pass, we do not verify later passes, such as the procedure to optimally configure the dataflow program on the actual hardware.
8. Related Work
As a compiler verification technique, the translation validation approach taken in our work is a less labor-intensive alternative to full compiler verification, such as CompCert (compcert, ). It was first proposed by Pnueli et al. (pnueli-tv, ), and there are numerous following works with novel translation validation or program equivalence checking techniques for various languages and compilers (necula-tv, ; kundu-tv, ; data-driven-eq, ; tate-eq-saturation, ). For LLVM-related works on translation validation, we have Alive (alive, ; alive2, ), which performs bounded translation validation for internal optimizations in LLVM. LLVM-MD (llvm-md, ; denotational-tv, ) is another translation validation tool for LLVM optimizations. LLVM-MD translates the LLVM code into an intermediate representation called synchronous value graphs (SVG) (denotational-tv, ), which is a similar representation to a dataflow program. The difference is that the semantics of SVG is synchronous, and they also trust the compilation from LLVM to SVG, whereas in our case, such translation is exactly what we are trying to validate. Our formulation of cut-bisimulation in our bisimulation check is an adaptation from the work on Keq (keq, ), which aims to be a more general program equivalence checker parametric in the operational semantics of input/output languages.
Our use of linear permission tokens is a dynamic variant of the fractional permissions proposed by Boyland (fractional-perm, ) and used in some variations of concurrent separation logic (concurrent-separation-logic, ). In contrast to separation logic frameworks such as Iris (iris, ), which are embedded in interactive proof assistants, our permission tokens are synthesized and checked fully automatically.
There are also verification tools for classical dataflow languages such as Lustre (lustre, ; lustre-verified-compiler, ; lustre-smt-verification, ) and Esterel (esterel-verification, ). These languages have a synchronous semantics and do not interact with globally accessible memory, whereas dataflow programs emitted by CGRA compilers are highly asynchronous and make pervasive use of the global heap.
In the context of term rewriting systems (term-rewriting, ), the standard approach for confluence checking is the use of critical pairs (knuth-critical-pairs, ; duran-confluence, ), which proves ground confluence in the case when the rewriting system is terminating. This algorithm is implemented in various systems, such as the Church-Rosser checker in Maude (maude, ). However, this technique is difficult to apply in our case, since the semantics of a dataflow program is neither terminating nor, in general, ground confluent.
9. Conclusion
In this work, we develop a translation validation technique for asynchronous dataflow programs with an application to the RipTide CGRA architecture. Our technique first checks the bisimulation of the input imperative program and the output dataflow program on a canonical schedule, and then we generalize the results to all possible schedules via linear permission tokens. Our work is an initial step towards integrating high assurance into new architectures from the start, rather than retrofitting them decades later.